A beam of protons, each of which has mass m, and charge qe moving with a velocity V enters perpendicularly a uniform magnetic field with the magnitude B= 2uT directed into plane of the page. After the beam exits the magnetic field, leaving the field in a direction perpendicular to its original direction, the beam enters a uniform electric field with the intensity E = 2.10-8 V/m directed downward existing above a distance d=10 m. The beam travels a distance of 6 cm in the magnetic field. After the proton beam exits the magnetic field, what will be passing time till the proton beam stops in the electric field? (Neglect the effect of gravitation).


1
Expert's answer
2021-08-31T08:58:09-0400


Since proton exists the magnetic field after travelling a quater circle, so the length of arc will traveled will be,

"L= \frac{2 \pi r}{4} \\

L= \frac{\pi r}{2}"

As per the question this length is equal to 6 cm.

"\frac{\pi r}{2}= 6 \times 10^{-2} \;m \\

\frac{3 r}{2} = 6 \times 10^{-2} \;m \\

r = 4 \times 10^{-2} \;m"

Radius of circle under went is equal to 4 cm. (π= 3)

Since Lorentz Magnetic force provides the necessary centripetal force,

"qvB = \frac{mv^2}{r} \\

r = \frac{mv}{qB}"

As seen from the figure the point of entry and exit for proton is equal to radius of circle under went.

"v=\frac{qBr}{m} \\

v = \frac{(1.6 \times 10^{-19})(2 \times 10^{-6}) (4 \times 10^{-2})}{1.67 \times 10^{-27}} \\

v = 7.66 \; m/s"

The proton enters and exits the region of magnetic field at 7.66 m/s as Lorentz force only produces change in direction.

The proton then moves 10 m in the region of no field, hence its kinetic energy hence velocity remains same, i.e proton enters electric field with a velocity of 7.66 m/s.

Once it enters the Electric field a Electric force, F = qE, opposes its motion hence,

"m_p\vec{a}= -q \vec{E}"

retardation produced,

"\vec{a} = -\frac{q \vec{E}}{m_p} \\

\vec{a} = - \frac{(1.6 \times 10^{-19})(2 \times 10^{-8})}{1.67 \times 10^{-27}} \\

\vec{a}= -1.92\; m/s"

Let S be the distance within the Electric field after which the proton comes to momentary rest.

By equation of motion,

"v^2=u^2 + 2aS \\

0^2 = u^2+ 2aS \\

0 = u^2 + 2(-1.92)S \\

u^2 = 3.84S \\

S=\frac{u^2}{3.84} \\

S = \frac {7.66^2}{3.92} \\

S = 14.97 \; m"

Time can be calculated using equation,

"v = u + at \\

0=u+ at \\

t = \frac{ u}{a} \\

t = \frac{7.66}{3.92} \\

t = 1.95 \; s"

Answer: 1.95 s